### v1

# The hull variation problem for projective Reed-Muller codes and quantum error-correcting codes

Abstract.
Long quantum codes using projective Reed-Muller codes are constructed. We obtain asymmetric and symmetric quantum codes by using the CSS construction and the Hermitian construction, respectively. Quantum codes obtained from projective Reed-Muller codes usually require entanglement assistance, but we show that sometimes we can avoid this requirement by considering monomially equivalent codes. Moreover, we also provide some constructions of quantum codes from subfield subcodes of projective Reed-Muller codes.
Key words and phrases:
Projective Reed-Muller codes, quantum codes, subfield subcodes, Hermitian product, hull
2020 Mathematics Subject Classification:
Primary: 81P70. Secondary: 94B05, 13P25
This work was supported in part by the following grants: Grant TED2021-130358B-I00 funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”, by Grant PID2022- 138906NB-C21 funded by MCIN/AEI/ 10.13039/501100011033 and by ERDF “A way of making Europe”, QCAYLE project funded by MCIN, the European Union NextGenerationEU (PRTR C17.I1) and Junta de Castilla y León and FPU20/01311 funded by the Spanish Ministry of Universities.
1. Introduction
Self-orthogonal linear codes with respect to the Euclidean (resp. Hermitian) inner product have been widely used for constructing quantum error-correcting codes (QECCs) using the CSS construction (resp. Hermitian construction) [6, 24, 16]. By using entanglement, it is possible to remove the self-orthogonality restriction [5]. Thus, we can use any linear code to construct entanglement-assisted quantum error-correcting codes (EAQECCs). For the CSS construction, one can consider two codes $C_{1}$ and $C_{2}$, and the minimum number required of maximally entangled quantum states is $c:=\dim C_{1}-\dim C_{1}\cap C_{2}^{\perp}$. Therefore, the dimension of the relative hull of $C_{1}$ with respect to $C_{2}$, which is defined in [1] as $\text{Hull}_{C_{2}}(C_{1}):=C_{1}\cap C_{2}^{\perp}$, determines the parameter $c$. For the Hermitian construction, we only use one code $C$, and the parameter $c$ is given by $\dim C-\dim C\cap C^{\perp_{h}}$, where $C^{\perp_{h}}$ is the Hermitian dual of $C$. The Hermitian hull of $C$ is thus defined as $\text{Hull}^{H}(C):=C\cap C^{\perp_{h}}$.
In this paper, we are interested in constructing QECCs with projective Reed-Muller codes, which are a family of evaluation codes obtained by evaluating homogeneous polynomials of a given degree at the projective space [17, 23]. In [17] it is shown that these codes can outperform affine Reed-Muller codes in some instances. In general, one needs to require entanglement assistance when working with evaluation codes over the projective space because the resulting families of codes are not necessarily nested. EAQECCs from projective Reed-Muller codes have been studied in [20] for the case of the projective plane, and in [22] for the few cases in which one has the nested condition.
To avoid requiring entanglement assistance, we can use the results from [1] and [7], which give conditions to increase the dimension of the relative hull and the Hermitian hull, respectively, using monomially equivalent codes. More propagation rules can be found in [18]. If we increase the dimension of the relative hull (respectively, Hermitian hull) to $\dim C_{1}$ (respectively, $\dim C$), then we obtain $c=0$ and the corresponding QECC does not require entanglement assistance. In this way, we find conditions to get unassisted QECC by using projective Reed-Muller codes, obtaining asymmetric and symmetric quantum codes with good parameters. One of the benefits of using projective Reed-Muller codes is that we can obtain long codes over a small finite field size by increasing the number of variables. Another approach to achieve that is the following: given $C\subset{\mathbb{F}}_{q^{s}}^{n}$, we consider
$C_{q}:=C\cap\mathbb{F}_{q}^{n},$
its subfield subcode with respect to the extension $\mathbb{F}_{q^{s}}/\mathbb{F}_{q}$. This technique has been used to construct quantum codes from projective Reed-Solomon codes [11], and has also been applied to projective Reed-Muller codes [12, 21]. As monomially equivalent codes do not have monomially equivalent subfield subcodes in general, using the techniques from [1] and [7] for projective Reed-Muller codes and then taking the subfield subcodes is not a very fruitful strategy. Nevertheless, we can still find conditions for the subfield subcodes of projective Reed-Muller codes to be nested, and therefore we can construct long QECCs over small finite fields using them.
2. Preliminaries
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements. We denote by ${\mathbb{P}^{m}}$ the projective space over $\mathbb{F}_{q}$. We choose for ${\mathbb{P}^{m}}$ the standard representatives, i.e., the representatives whose leftmost coordinate is equal to 1. If we regard the standard representatives as points in the affine space $\mathbb{A}^{m+1}$, we obtain the set $P^{m}:=\{Q_{1},\dots,Q_{n}\}\subset\mathbb{A}^{m+1}$, where $n=\left\lvert P^{m}\right\rvert=\frac{q^{m+1}-1}{q-1}$.
Let $S=\mathbb{F}_{q}[x_{0},\dots,x_{m}]$. Given $d$ a positive integer, we denote $S_{d}\subset S$ the set of homogeneous polynomials of degree $d$. We define the evaluation map
$\operatorname{ev}:S_{d}\rightarrow\mathbb{F}_{q}^{n},\>\>f\mapsto\left(f(Q_{1}),\dots,f(Q_{n})\right).$
The image of $\operatorname{ev}$ is the projective Reed-Muller code of degree $d$, denoted by $\operatorname{PRM}_{d}(q,m)$, or $\operatorname{PRM}_{d}(m)$ if there is no confusion about the field. For the minimum distance of a code $C\subset\mathbb{F}_{q}^{n}$, we use the notation $\operatorname{wt}(C)$. We have the following results about the parameters of projective Reed-Muller codes and their duality from [23].
Theorem 2.1.
The projective Reed-Muller code $\operatorname{PRM}_{d}(q,m)$, $1\leq d\leq m(q-1)$, is an $[n,k]$-code with
$n=\frac{q^{m+1}-1}{q-1},$
$k=\sum_{t\equiv d\bmod q-1,0<t\leq d}\left(\sum_{j=0}^{m+1}(-1)^{j}\binom{m+1}{j}\binom{t-jq+m}{t-jq}\right).$
For the minimum distance, we have
$\operatorname{wt}(\operatorname{PRM}_{d}(q,m))=(q-\ell)q^{m-r-1},\text{ where }\;d-1=r(q-1)+\ell,\;0\leq\ell<q-1.$
Theorem 2.2.
Let $1\leq d\leq m(q-1)$ and let $d^{\perp}=m(q-1)-d$. Then
[…]
With respect to affine Reed-Muller codes, we will denote them by $\operatorname{RM}_{d}(q,m)$, or by $\operatorname{RM}_{d}(m)$ if there is no confusion about the field. We recall now the analogous results for affine Reed-Muller codes [8, 15].
Theorem 2.3.
The Reed-Muller code $\operatorname{RM}_{d}(q,m)$, $1\leq d\leq m(q-1)$, is an $[n,k]$-code with
$n=q^{m},$
$k=\sum_{t=0}^{d}\sum_{j=0}^{m}(-1)^{j}\binom{m}{j}\binom{t-jq+m-1}{t-jq}.$
For the minimum distance, we have
$\operatorname{wt}(\operatorname{RM}_{d}(q,m))=(q-\ell)q^{m-r-1},\text{ where }\;d=r(q-1)+\ell,\;0\leq\ell<q-1.$
Theorem 2.4.
Let $1\leq d\leq m(q-1)$. Then
$\operatorname{RM}_{d}^{\perp}(q,m)=\operatorname{RM}_{m(q-1)-d-1}(q,m).$
For comparisons between the projective case and the affine case, it is important to note that $\operatorname{wt}(\operatorname{PRM}_{d}(q,m))=\operatorname{wt}(\operatorname{RM}_{d-1}(q,m))$.
Regarding quantum codes, we will consider the CSS construction and the Hermitian construction. For the CSS construction, we are considering asymmetric QECCs. In quantum error-correction we may consider two different types of errors, phase-shift errors and qudit-flip errors. As the likelihood of occurrence of each type of error is different [14], it is desirable to have different error correction capabilities for each type of error, which is what asymmetric QECCs accomplish. Asymmetric QECCs have two minimum distances, $\delta_{z}$ and $\delta_{x}$, meaning that this code can correct up to $\lfloor(\delta_{z}-1)/2\rfloor$ phase-shift errors and $\lfloor(\delta_{x}-1)/2\rfloor$ qudit-flip errors. We state now the CSS construction for asymmetric EAQECCs [10].
Theorem 2.5 ((CSS Construction)).
Let $C_{i}\subset\mathbb{F}_{q}^{n}$ be linear codes of dimension $k_{i}$, for $i=1,2$. Then, there is an asymmetric EAQECC with parameters $[[n,\kappa,\delta_{z}/\delta_{x};c]]_{q}$, where
$c$ $=k_{1}-\dim(C_{1}\cap C_{2}^{\perp}),\;\kappa=n-(k_{1}+k_{2})+c,$
$\delta_{z}=\operatorname{wt}$ $\left(C_{1}^{\perp}\setminus\left(C_{1}^{\perp}\cap C_{2}\right)\right)\text{ and }\;\delta_{x}=\operatorname{wt}\left(C_{2}^{\perp}\setminus\left(C_{2}^{\perp}\cap C_{1}\right)\right).$
As we stated in the introduction, we are interested in constructing QECCs, withouth entanglement assistance, which corresponds to pairs of codes $C_{1},C_{2}$ such that $\dim\text{Hull}_{C_{2}}(C_{1})=\dim C_{1}$, equivalently, $C_{1}\subset C_{2}^{\perp}$.
For the Hermitian construction, we need to consider codes defined over ${\mathbb{F}}_{q^{2}}^{n}$ and the Hermitian product. For two vectors $v,w\in{\mathbb{F}}_{q^{2}}^{n}$, their Hermitian product is
$v\cdot_{h}w:=\sum_{i=1}^{n}v_{i}w_{i}^{q}.$
The Hermitian dual of a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$ is defined as $C^{\perp_{h}}:=\{v\in{\mathbb{F}}_{q^{2}}^{n}\mid v\cdot_{h}w=0,\;\forall\;w\in C\}$.
Remark 2.6.
For a code $C\subset{\mathbb{F}}_{q^{2}}^{n}$, we have that $C^{\perp_{h}}=(C^{\perp})^{q}$, where we consider the component wise power of $q$. This implies that the Euclidean dual and the Hermitian dual have the same parameters.
We can state now the Hermitian construction for EAQECCs [10].
Theorem 2.7 ((Hermitian construction)).
Let $C\subset{\mathbb{F}}_{q^{2}}^{n}$ be a linear code of dimension $k$ and $C^{\perp_{h}}$ its Hermitian dual. Then, there is an EAQECC with parameters $[[n,\kappa,\delta;c]]_{q}$, where
$c=k-\dim(C\cap C^{\perp_{h}}),\;\kappa=n-2k+c,\;\text{ and }\;\delta=\operatorname{wt}(C^{\perp_{h}}\setminus(C\cap C^{\perp_{h}})).$
From the Hermitian construction for EAQECCs we can only obtain symmetric EAQECCs, that is, EAQECCs with $\delta_{z}=\delta_{x}=\delta$. As we stated before, we are interested in codes such that $\dim\text{Hull}^{H}(C)=\dim C$, equivalently, $C\subset C^{\perp_{h}}$.
For asymmetric QECCs (EAQECCs with $c=0$), we will use the notation $[[n,\kappa,\delta_{z}/\delta_{x}]]_{q}$ for their parameters. Analogously, for symmetric QECCs we will use $[[n,\kappa,\delta]]_{q}$. In what follows, we will assume that the QECCs we obtain are pure (if they were impure, then the minimum distance that we give would be a lower bound for the true minimum distance). This corresponds to assuming $\delta_{z}=\operatorname{wt}(C_{1}^{\perp})$ and $\delta_{x}=\operatorname{wt}(C_{2}^{\perp})$ ($\delta=\operatorname{wt}(C^{\perp_{h}})$ for the Hermitian construction).
3. QECCs from the CSS construction
In this section we focus on the QECCs obtained using the CSS Construction 2.5 with projective Reed-Muller codes. We do this by increasing the size of the relative hull via monomially equivalent codes. We give some preliminaries first. Given two vectors $v,u\in\mathbb{F}_{q}^{n}$, we use the following notation:
$u\star v:=(u_{1}v_{1},\dots,u_{n}v_{n})\in\mathbb{F}_{q}^{n}.$
For two codes $C_{1},C_{2}$ in $\mathbb{F}_{q}^{n}$, we consider
$C_{1}\star C_{2}:=\langle u_{1}\star u_{2}\mid u_{1}\in C_{1},\;u_{2}\in C_{2}\rangle.$
We say that $C_{1}$ is monomially equivalent to $C_{2}$ if there is some vector $v\in\mathbb{F}_{q}^{n}$ with Hamming weight $n$ such that $C_{1}=\langle v\rangle\star C_{2}$. It is clear that monomially equivalent codes have the same basic parameters. The following result from [1] allows us to increase the size of the relative hull in some cases.
Theorem 3.1.
For $i=1,2$, let $C_{i}$ be $[n,k_{i}]_{q}$ codes with $q>2$ and $k_{1}+k_{2}<n$. If there is a codeword of Hamming weight $n$ in $(C_{1}\star C_{2})^{\perp}$, then for any $\dim\operatorname{Hull}_{C_{2}}(C_{1})\leq l\leq k_{1}$, there exists a monomially equivalent code $C_{2,l}$ to $C_{2}$ such that $\dim\operatorname{Hull}_{C_{2,l}}(C_{1})=l$.
The idea is to use Theorem 3.1 with projective Reed-Muller codes. The following lemma is key for the subsequent results.
Lemma 3.2.
Let $\gamma$ be a non-negative integer, and $x^{\gamma}\in\mathbb{F}_{q}[x]$. We have the following:
$\sum_{z\in{\mathbb{F}}_{q}}x^{\gamma}(z)=\begin{cases}0&\text{ if }\gamma=0\text{ or }\gamma>0\text{ and }\gamma\not\equiv 0\bmod(q-1),\\ -1&\text{ if }\gamma>0\text{ and }\gamma\equiv 0\bmod(q-1).\end{cases}$
Proof.
Let $\xi\in{\mathbb{F}}_{q}$ be a primitive element. Then $\mathbb{F}_{q}=\{\xi^{0},\xi^{1},\dots,\xi^{N-2}\}\cup\{0\}$. If $\gamma=0$, $x^{\gamma}=1$, and the sum is equal to $\left\lvert\mathbb{F}_{q}\right\rvert=q=0$ in $\mathbb{F}_{q}$. If $\gamma>0$ and $\gamma\equiv 0\bmod(q-1)$, then $x^{\gamma}(z)=1$ for all $z\in\mathbb{F}_{q}^{*}$, and $\sum_{z\in\mathbb{F}_{q}}x^{\gamma}(z)=q-1=-1$. Finally, if $\gamma>0$ and $\gamma\not\equiv 0\bmod(q-1)$, we have
$\sum_{z\in\mathbb{F}_{q}}x^{\gamma}(z)=\sum_{i=0}^{q-2}(\xi^{i})^{\gamma}=\frac{\xi^{\gamma(q-1)}-1}{\xi^{\gamma}-1}=0.$
Remark 3.3.
When working over the affine space $\mathbb{A}^{\ell}$, $1\leq\ell\leq m$, if we consider $x_{1}^{\alpha_{1}}\cdots x_{\ell}^{\alpha_{\ell}}\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]$, we have
$\sum_{Q\in\mathbb{A}^{\ell}}x_{1}^{\alpha_{1}}\cdots x_{\ell}^{\alpha_{\ell}}(Q)=\left(\sum_{z\in\mathbb{F}_{q}}x_{1}^{\alpha_{1}}(z)\right)\cdots\left(\sum_{z\in\mathbb{F}_{q}}x_{\ell}^{\alpha_{\ell}}(z)\right).$
Thus, we can use Lemma 3.2 in order to obtain the result of this sum. In particular, if we have $\alpha_{i}<q-1$ for some $1\leq i\leq\ell$, this sum is equal to 0.
Lemma 3.4.
Let $21\leq d<q-$ 2. Then there is a vector of Hamming weight $n=\frac{q^{m+1}-1}{q-1}$ in $\operatorname{PRM}_{d}^{\perp}(m)$.
Proof.
Let $t$ be a monic polynomial of degree $q-1-d$ such that $t(z)\neq 0$ for every $z\in\mathbb{F}_{q}$. For example, we can choose $t$ as a monic irreducible polynomial in $\mathbb{F}_{q}[x]$. We consider the vector $v=(v_{Q})_{Q\in P^{m}}\in\mathbb{F}_{q}^{n}$ defined in the following way: if $Q=(0,0,\dots,0,1,z)$, for some $z\in\mathbb{F}_{q}$, we define $v_{Q}=t(z)$, and $v_{Q}=1$ otherwise. For $0\leq i\leq q-1-d$, let $t_{i}$ be the coefficient of $x^{i}$ in $t$. We consider the following decomposition of $P^{m}=(\{1\}\times{\mathbb{F}}_{q}^{m})\cup(\{0\}\times\{1\}\times{\mathbb{F}}_{q}^{m-1})\cup\cdots\cup\{(0,0,\dots,0,1)\}=B_{m}\cup B_{m-1}\cup\cdots\cup B_{0}$. Let $x_{0}^{\alpha_{0}}x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}}=x^{\alpha}\in\mathbb{F}_{q}[x_{0},\dots,x_{m}]_{d}$. Then we have
$v\cdot\operatorname{ev}(x^{\alpha})=\sum_{Q\in B_{m}}v_{Q}x^{\alpha}(Q)+\sum_{Q\in B_{m-1}}v_{Q}x^{\alpha}(Q)+\cdots+\sum_{Q\in B_{0}}v_{Q}x^{\alpha}(Q).$
For $2\leq i\leq m$, we have $v_{Q}=1$ and
$\sum_{Q\in B_{i}}v_{Q}x^{\alpha}(Q)=\sum_{Q\in B_{i}}x^{\alpha}(Q)=0$
by Remark 3.3 because $\alpha_{j}<q-1$ for $0\leq j\leq m$.
We study the sum over $B_{1}$ now. If $\alpha_{j}=0$ for every $20\leq j\leq m-$ 2, we have
(1) $\sum_{Q=[0:0:\cdots:1:z]\in B_{1}}v_{Q}x^{\alpha}(Q)=\sum_{z\in\mathbb{F}_{q}}t(z)x_{m}^{\alpha_{m}}(z)=\sum_{l=0}^{q-1-d}t_{l}\sum_{z\in\mathbb{F}_{q}}x_{m}^{\alpha_{m}+l}(z).$
Taking into account that $\alpha_{m}\leq d$, we have $\alpha_{m}+l\leq q-1$, and we can only have the equality for some $l\in\{0,\dots,q-1-d\}$ if $\alpha_{m}=d$. Therefore, by Lemma 3.2, this sum is equal to 0 unless $\alpha_{m}=d$, in which case it is equal to $-t_{q-1-d}=-1$ ($t$ is monic).
On the other hand, if we have $\alpha_{j}>0$ for some $20\leq j\leq m-$ 2, then the sum over $B_{1}$ in (1) is clearly equal to 0 because all the addends are equal to 0. For the sum over $B_{0}$, it is clear that this sum is equal to 0 unless $\alpha_{m}=d$, in which case it is equal to 1. Hence, if $\alpha_{m}\neq d$, all the sums are equal to 0 and we have $v\cdot\operatorname{ev}(x^{\alpha})=0$, and if $\alpha_{m}=d$, all the sums are 0 besides the sums corresponding to $B_{1}$ and $B_{0}$, which are equal to $-1$ and $1$, respectively. Thus, if $\alpha_{m}=d$ we also have $v\cdot\operatorname{ev}(x_{m}^{d})=0$. This implies that $v\in\operatorname{PRM}_{d}^{\perp}(m)$, and the fact that $v$ has Hamming weight $n$ follows from its definition. ∎
Note that, for $d\equiv 0\bmod q-1$, $d\leq\lfloor m(q-1)/2\rfloor$, we have $(1,\dots,1)$ in $\operatorname{PRM}_{d}^{\perp}(m)$ by Theorem 2.2, which is a vector of Hamming weight $n$, and we also know that the corresponding projective Reed-Muller code is contained in its dual [22, Cor. 10.3]. As a consequence of Theorem 3.1 and Lemma 3.4, we obtain the following.
Proposition 3.5.
Let $q>2$ and let $21\leq d_{1}\leq d_{2}<q-$ 2. Let $C_{i}=\operatorname{PRM}_{d_{i}}(m)$, for $i=1,2$. If $d_{1}+d_{2}<q-2$, then, for any $\dim C_{1}\cap C_{2}^{\perp}\leq l\leq\dim C_{1}$, there is a code $C_{2,l}$ monomially equivalent to $C_{2}$ such that
$\dim C_{1}\cap C_{2,l}^{\perp}=l.$
Proof.
We have that $C_{1}\star C_{2}=\operatorname{PRM}_{d_{1}+d_{2}}(m)$, with $d_{1}+d_{2}<q-2$. By Lemma 3.4, we have a vector of Hamming weight $n$ in $(C_{1}\star C_{2})^{\perp}$. Using Theorem 3.1, we obtain the result. ∎
We can use Proposition 3.5 and Theorem 2.5 to construct asymmetric and symmetric QECCs.
Theorem 3.6.
Let $21\leq d_{1}\leq d_{2}<q-$ 2 such that $d_{1}+d_{2}<q-2$. Then we can construct a quantum code with parameters $[[n,\kappa,\delta_{z}/\delta_{x}]]_{q}$, where $n=\frac{q^{m+1}-1}{q-1}$, $\kappa=n-(\dim\operatorname{PRM}_{d_{1}}(m)+\dim\operatorname{PRM}_{d_{2}}(m))$, $\delta_{z}=\operatorname{wt}(\operatorname{PRM}_{d_{2}}^{\perp}(m))$ and $\delta_{x}=\operatorname{wt}(\operatorname{PRM}_{d_{1}}^{\perp}(m))$.
The previous result, with $m=2$, removes the entanglement requirement for some of the codes arising from [20].
Remark 3.7.
For $21\leq d_{1}\leq d_{2}<q-$ 2 such that $d_{1}+d_{2}<q-2$, by Theorem 2.3 and Theorem 2.1 we have $\dim\operatorname{RM}_{d_{i}}(m)=\sum_{t=0}^{d_{i}}\binom{t+m-1}{t}=\binom{d_{i}+m}{m}=\dim\operatorname{PRM}_{d_{i}}(m)$, for $i=1,2$. Moreover, we have that $\operatorname{wt}(\operatorname{PRM}_{d_{i}}^{\perp}(m))=\operatorname{wt}(\operatorname{PRM}_{d_{i}^{\perp}}(m))=\operatorname{wt}(\operatorname{RM}_{d_{i}^{\perp}-1}(m))=\operatorname{wt}(\operatorname{RM}_{d_{i}}^{\perp}(m))$, where $d_{i}^{\perp}=m(q-1)-d_{i}$, for $i=1,2$. Therefore, using Theorem 2.5 with $C_{i}=\operatorname{RM}_{d_{i}}(m)$ for $i=1,2$, we get an asymmetric QECC with parameters
$[[q^{m},q^{m}-(\dim\operatorname{RM}_{d_{1}}(m)+\dim\operatorname{RM}_{d_{2}}(m)),\operatorname{wt}(\operatorname{RM}_{d_{2}}^{\perp}(m))/\operatorname{wt}(\operatorname{RM}_{d_{1}}^{\perp}(m))]]_{q}.$
On the other hand, using Theorem 3.6 we get an asymmetric QECC with parameters
$[[q^{m}+\Delta,q^{m}+\Delta-(\dim\operatorname{PRM}_{d_{1}}(m)+\dim\operatorname{PRM}_{d_{2}}(m)),\operatorname{wt}(\operatorname{PRM}_{d_{2}}^{\perp}(m))/\operatorname{wt}(\operatorname{PRM}_{d_{1}}^{\perp}(m))]]_{q},$
where $\Delta=\frac{q^{m}-1}{q-1}$. Taking into account the previous discussion, we see that both asymmetric QECCs have the same minimum distances $\delta_{z}$ and $\delta_{x}$, but the code obtained using Theorem 3.6 has gained $\Delta$ units in length and dimension, which increases the code rate and decreases the relative minimum distances. If we consider $\frac{\kappa+\delta_{z}}{n}$ (resp. $\frac{\kappa+\delta_{x}}{n}$) as a measure of how good a code is in terms of transmission rate and phase-shift error-correction capability (resp. qudit-flip error-correction capability), we see that in this case we obtain asymmetric QECCs with better performance using projective Reed-Muller codes (Theorem 3.6) than the asymmetric QECCs obtained using affine Reed-Muller codes.
As there are not many tables to compare asymmetric QECCs, the following quantum Gilbert-Varshamov bound from [19] is useful to check that the codes we obtain are good.
Theorem 3.8.
Assume the existence of integers $n\geq 1$, $1\leq l<n+c$, $\delta_{x}\geq 1$, $\delta_{z}\geq 1$, $20\leq c\leq l/$ 2 such that
$\frac{q^{2n-l}-q^{l-2c}}{q^{2n}-1}\left(\sum_{i=0}^{\delta_{x}-1}\binom{n}{i}(q-1)^{i}\sum_{j=0}^{\delta_{z}-1}\binom{n}{i}(q-1)^{j}-1\right)<1.$
Then there is an EAQECC with parameters $[[n,n-l+c,\delta_{z}/\delta_{x};c]]_{q}$.
Given a quantum Gilbert-Varshamov bound, like that one from Theorem 3.8, we will say that a code $C$ surpasses it if the bound does not guarantee the existence of a code with the parameters of $C$.
Example 3.9.
Let $q=8$, $m=2$, $d_{1}=1$ and $d_{2}=4$. We have $d_{1}+d_{2}=5<q-2$, and we can apply Theorem 3.6 to obtain a QECC with parameters $[[73,55,6/3]]_{8}$. This code surpasses the Gilbert-Varshamov bound from Theorem 3.8.
In some cases it is possible to obtain nested pairs of codes from subfield subcodes of projective Reed-Muller codes, giving rise to QECCs, as we show in the next result.
Proposition 3.10.
Let $d_{1},d_{2}$ such that $d_{1}+d_{2}=\lambda(q^{s}-1)$, with $1\leq\lambda\leq m$. Then we have $(\operatorname{PRM}_{d_{1}}(q^{s},m))_{q}\subset((\operatorname{PRM}_{d_{2}}(q^{s},m))_{q})^{\perp}$, and we can construct a QECC with parameters $[[n,\kappa,\delta]]_{q}$, where $n=\frac{q^{s(m+1)}-1}{q^{s}-1}$, $\kappa=n-\dim(\operatorname{PRM}_{d_{1}}(q^{s},m))_{q}-\dim(\operatorname{PRM}_{d_{2}}(q^{s},m))_{q}$, $\delta_{z}\geq\operatorname{wt}(((\operatorname{PRM}_{d_{2}}(q^{s},m))_{q})^{\perp})$ and $\delta_{x}\geq\operatorname{wt}(((\operatorname{PRM}_{d_{1}}(q^{s},m))_{q})^{\perp})$.
Proof.
By hypothesis we have $d_{1}\equiv d_{2}^{\perp}\bmod q^{s}-1$ and $d_{1}\leq d_{2}^{\perp}$. By [22, Lem. 10.7], $\operatorname{PRM}_{d_{1}}(q^{s},m)\subset\operatorname{PRM}_{d_{2}}^{\perp}(q^{s},m)$. This implies
$(\operatorname{PRM}_{d_{1}}(q^{s},m))_{q}\subset(\operatorname{PRM}_{d_{2}}^{\perp}(q^{s},m))_{q}\subset((\operatorname{PRM}_{d_{2}}(q^{s},m))_{q})^{\perp}.$
The parameters follow from Theorem 2.5. ∎
Remark 3.11.
As we do not know in general $\operatorname{wt}(((\operatorname{PRM}_{d_{i}}(q^{s},m))_{q})^{\perp})$, for $i=1,2$, if we want to use the previous result we have to compute this minimum distance with Magma [4]. Moreover, if we want to have minimum distances greater than 1, we have to ensure that $(\operatorname{PRM}_{d_{i}}(q^{s},m))_{q}$, for $i=1,2$, are non degenerate. The results from [21] give us some cases in which those subfield subcodes are non degenerate, and in those cases we also have formulas for the dimension of the subfield subcodes. For $m=2$, we can also use the results from [12].
Example 3.12.
Let $m=2=s$. From [21] (or [12]), if $d_{i}\equiv 0\bmod(q^{s}-1)/(q-1)$, then $(\operatorname{PRM}_{d_{i}}(q^{s},2))_{q}$ is non degenerate, for $i=1,2$. If we set $q=2$, we obtain a code with parameters $[[73,19,9/9]]_{2}$ with $d_{1}=d_{2}=7$. For $q=3$, we obtain the parameters $[[91,73,4/4]]_{3}$ with $d_{1}=d_{2}=4$, and $[[91,12,36/4]]_{2}$ with $d_{1}=4$, $d_{2}=12$. All of them surpass the Gilbert-Varshamov bound from Theorem 3.8.
4. QECCs from the Hermitian construction
In the Hermitian case, we can also obtain quantum codes with projective Reed-Muller codes without entanglement assistance. Recall that in this case we have to consider codes over ${\mathbb{F}}_{q^{2}}$, and the Hermitian product. The following result from [7, Thm. 2.1] is analogous to Theorem 3.1 for the Hermitian product.
Theorem 4.1.
Let $C\subset{\mathbb{F}}_{q^{2}}$ be a linear code. If there is a vector $v\in((C\star C^{q})_{q})^{\perp}$ with $\operatorname{wt}(v)=n$, then $\langle v\rangle\star C\subset(\langle v\rangle\star C)^{\perp_{h}}$, i.e., $\langle v\rangle\star C$ is self-orthogonal with respect to the Hermitian product.
By using the previous result we can easily find some projective Reed-Muller which are self-orthogonal with respect to the Hermitian product.
Proposition 4.2.
Let $d=\lambda(q-1)$ with $1\leq\lambda\leq m$. Then $\operatorname{PRM}_{d}(q^{2},m)\subset\operatorname{PRM}_{d}^{\perp_{h}}(q^{2},m)$. As a consequence, if $d\equiv 0\bmod q-1$ and $d\not\equiv 0\bmod q^{2}-1$, we can construct a QECC with parameters $[[n,\kappa,\delta]]_{q}$, where $n=\frac{q^{2(m+1)}-1}{q^{2}-1}$, $\kappa=n-2(\dim\operatorname{PRM}_{d}(q^{2},m))$ and $\delta\geq\operatorname{wt}(\operatorname{PRM}_{d}^{\perp}(q^{2},m))$.
Proof.
We have
$((\operatorname{PRM}_{d}(q^{2},m)\star\operatorname{PRM}_{d}(q^{2},m)^{q})_{q})^{\perp}$ $\supset((\operatorname{PRM}_{d}(q^{2},m)\star\operatorname{PRM}_{d}(q^{2},m)^{q})^{\perp})_{q}$
$\supset(\operatorname{PRM}_{d+qd}^{\perp}(q^{2},m))_{q}.$
By Theorem 2.2, if we have $d(q+1)\equiv 0\bmod q^{2}-1$, with $d(q+1)\leq m(q^{2}-1)$, then $(1,\dots,1)\in(\operatorname{PRM}_{d+qd}^{\perp}(q^{2},m))_{q}$ and we can apply Theorem 4.1 with $C=\operatorname{PRM}_{d}(q^{2},m)$. The parameters of the quantum code are deduced from Theorems 2.1, 2.2 and 2.7. ∎
The codes that we obtain from this construction (and, in general, in this section) are usually quite long. To check if this codes are good, we introduce now another quantum Gilbert-Varshamov bound from [9], which seems to be more difficult to surpass for symmetric QECCs than the bound from Theorem 3.8.
Theorem 4.3.
Suppose that $n>\kappa$, $\delta\geq 2$ and $n\equiv\kappa\bmod 2$. Then there exists a pure stabilizer quantum code $[[n,\kappa,\delta]]_{q}$, provided that
$\frac{q^{n-\kappa+2}-1}{q^{2}-1}>\sum_{i=1}^{\delta-1}(q^{2}-1)^{i-1}\binom{n}{i}.$
Example 4.4.
Let $q=2$ and $m=3$. For $d=q-1=1$ we can apply Proposition 4.2 to obtain a QECC with parameters $[[85,77,3]]_{2}$, which is optimal according to [13]. If we assume $m=4$ instead, we get a QECC with parameters $[[341,331,3]]_{2}$, which surpasses the quantum Gilbert-Varshamov bound from Theorem 4.3.
Consider now $q=4$ and $m=2$. For $d=q-1=3$ we obtain the parameters $[[273,253,5]]_{4}$, surpassing the quantum Gilbert-Varshamov bound from Theorem 4.3.
When $d\not\equiv 0\bmod q-1$, we can still reduce the dimension of the hull (which reduces the parameter $c$) via monomially equivalent codes in some cases. The idea of the proof of the following result is similar to the proof of [2, Thm. 4].
Theorem 4.5.
Let $21\leq d<q-$ 2. Then we can construct a QECC with parameters $[[n,\kappa,\delta]]_{q}$, where $n=\frac{q^{2(m+1)}-1}{q^{2}-1}$, $\kappa=n-2(\dim\operatorname{PRM}_{d}(q^{2},m))$ and $\delta\geq\operatorname{wt}(\operatorname{PRM}_{d^{\perp}}(q^{2},m))$.
Proof.
We will use Theorem 4.1 with $C=\operatorname{PRM}_{d}(q^{2},m)$. First, we find $w\in((C\star C^{q})^{\perp})_{q}\subset((C\star C^{q})_{q})^{\perp}$ with $\operatorname{wt}(v)=n$. Let $t$ be a monic polynomial with coefficients in ${\mathbb{F}}_{q^{2}}$ of degree $q-1-d$ such that $t(z)\neq 0$ for every $z\in{\mathbb{F}}_{q^{2}}$, and $v\in{\mathbb{F}}_{q^{2}}^{n}$ as in Lemma 3.4. Then we consider $w:=v^{q+1}$, where the power is taken component wise. Clearly $w\in{\mathbb{F}}_{q}^{n}$, $\operatorname{wt}(w)=n$, and we will prove that $w\in(C\star C^{q})^{\perp}$. Let $x^{\alpha},x^{\beta}\in{\mathbb{F}}_{q^{2}}[x_{0},\dots,x_{m}]_{d}$. Using the decomposition $P^{m}=B_{m}\cup B_{m-1}\cup\cdots\cup B_{0}$ from the proof of Lemma 3.4 we have that
$w\cdot(\operatorname{ev}(x^{\alpha})\star\operatorname{ev}(x^{q\beta}))=\sum_{i=0}^{m}\sum_{Q\in B_{i}}v_{Q}^{q+1}x^{\alpha+q\beta}(Q).$
For $2\leq i\leq m$, we have $v_{Q}=1$ and
$\sum_{Q\in B_{i}}v_{Q}^{q+1}x^{\alpha+q\beta}(Q)=\sum_{Q\in B_{i}}x^{\alpha+q\beta}(Q)=0$
because $\alpha_{j}+q\beta_{j}<q-1+q(q-1)=q^{2}-1$ for $0\leq j\leq m$.
On the other hand, if $\alpha_{j}=0$ for $20\leq j\leq m-$ 2, we have
(2) $\sum_{Q=[0:0:\cdots:1:z]\in B_{1}}v_{Q}^{q+1}x^{\alpha+q\beta}(Q)=\sum_{z\in{\mathbb{F}}_{q^{2}}}t(z)^{q+1}z^{\alpha_{m}+q\beta_{m}}=\sum_{l=0}^{(q+1)(q-1-d)}r_{l}\sum_{z\in{\mathbb{F}}_{q^{2}}}z^{l+\alpha_{m}+q\beta_{m}},$
where $r_{l}$ is the coefficient of $x^{l}$ in $t^{q+1}$. Taking into account that $l+\alpha_{m}+q\beta_{m}\leq(q+1)(q-1-d)+d+qd=q^{2}-1$, we see that this sum is equal to 0 unless $\alpha_{m}=\beta_{m}=d$, in which case it is equal to $-r_{(q+1)(q-1-d)}=-1$ (because $t$ is monic). If we have $\alpha_{j}>0$ for some $20\leq j\leq m-$ 2, then all the addends in (2) are equal to 0. For the sum over $B_{0}$, it is clear that this sum is equal to 0 unless $\alpha_{m}=\beta_{m}=d$, in which case it is equal to 1.
Therefore, if $\alpha_{m}\neq d$ or $\beta_{m}\neq d$, all the sums are equal to 0 and we have $w\cdot(\operatorname{ev}(x^{\alpha})\star\operatorname{ev}(x^{q\beta}))=0$, and if $\alpha_{m}=\beta_{m}=d$, all the sums are 0 besides the sums corresponding to $B_{1}$ and $B_{0}$, which are equal to $-1$ and $1$, respectively. Thus, if $\alpha_{m}=\beta_{m}=d$ we also have $w\cdot(\operatorname{ev}(x^{\alpha})\star\operatorname{ev}(x^{q\beta}))=0$. This implies that $w\in(C\star C^{q})^{\perp}$. Therefore, $\langle w\rangle\star C\subset(\langle w\rangle\star C)^{\perp_{h}}$ by Theorem 4.1, and we finish the proof by applying the Hermitian construction from Theorem 2.7 to $\langle w\rangle\star C$. ∎
Similarly to what we stated in Remark 3.7, in this situation we can obtain codes with $\frac{q^{2m}-1}{q^{2}-1}$ extra length and dimension with respect to the affine case.
Example 4.6.
We consider $q=4$ and $m=2$. Therefore, we work over the field ${\mathbb{F}}_{4^{2}}$ and we obtain codes of length $\frac{q^{2(m+1)}-1}{q^{2}-1}=273$. For $d=1$ we can apply Theorem 4.5 to obtain a QECC with parameters $[[273,267,3]]_{4}$. This code improves the parameters of the code $[[273,265,3]]_{4}$ from [3] and surpasses the quantum Gilbert-Varshamov from Theorem 4.3.
For $q=5$, $m=2$ and $d=1,2,$ we obtain the parameters $[[651,645,3]]_{5}$ and $[[651,639,4]]_{5}$, respectively. The first one improves the parameters $[[651,642,3]]_{5}$ and $[[652,644,3]]_{5}$ obtained in [3], and the second one improves the parameters $[[651,636,4]]_{5}$ and $[[652,638,4]]_{5}$ from [3]. Both of them exceed the quantum Gilbert-Varshamov bound from Theorem 4.3.
Similarly to the previous section, we can also use subfield subcodes of projective Reed-Muller codes in some cases. For the Hermitian product, we consider projective Reed-Muller codes over ${\mathbb{F}}_{q^{2s}}$ such that their subfield subcodes with respect to the extension ${\mathbb{F}}_{q^{2s}}/{\mathbb{F}}_{q^{2}}$ are self-orthogonal with respect to the Hermitian product, which gives rise to QECCs over ${\mathbb{F}}_{q}$ using the Hermitian construction from Theorem 2.7.
Proposition 4.7.
Let $d=\lambda(q^{2s}-1)/(q+1)$ with $1\leq\lambda\leq m$. Then $(\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}}\subset((\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}})^{\perp_{h}}$. As a consequence, we can construct a QECC with parameters $[[n,\kappa,\delta]]_{q}$, where
$n=\frac{q^{2s(m+1)}-1}{q^{2s}-1},\;\kappa=n-2(\dim(\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}})\text{ and }\delta\geq\operatorname{wt}(((\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}})^{\perp}).$
Proof.
We have
$(((\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}}\star((\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}})^{q})_{q})^{\perp}\supset((\operatorname{PRM}_{d}(q^{2s},m)\star(\operatorname{PRM}_{d}(q^{2s},m))^{q})_{q})^{\perp}$
$\supset((\operatorname{PRM}_{d+qd}(q^{2s},m))_{q})^{\perp}\supset(\operatorname{PRM}_{d+qd}^{\perp}(q^{2s},m))_{q}.$
The rest of the proof follows as in the proof of Proposition 4.2, changing $q^{2}$ with $q^{2s}$. ∎
As we stated in Remark 3.11, we have to use Magma [4] to compute the minimum distance if we want to use the previous result, and we have to ensure that $(\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}}$ is non degenerate. In the next example we show a possible way to do it using the results from [21].
Example 4.8.
By [21, Cor. 4.1 and Cor 4.2], if $d\equiv 0\bmod(q^{2s}-1)/(q^{2}-1)$, then we have recursive formulas for the dimension of $(\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}}$, and we know that $(\operatorname{PRM}_{d}(q^{2s},m))_{q^{2}}$ is non degenerate (this is also seen recursively, using what we know for the case $m=1$ from [11]).
Let $q=2$ and $m\geq 2$. Then $(q^{2s}-1)/(q+1)=(q^{2s}-1)/(q^{2}-1)=(4^{s}-1)/3$. Hence, by Proposition 4.7, if we consider $d=\lambda(4^{s}-1)/3$, for some $1\leq\lambda\leq m$, then we have $(\operatorname{PRM}_{d}(4^{s},m))_{q^{2}}\subset((\operatorname{PRM}_{d}(4^{s},m))_{q^{2}})^{\perp_{h}}$, and by the previous discussion $((\operatorname{PRM}_{d}(4^{s},m))_{q^{2}})^{\perp_{h}}$ is non degenerate. For example, for $m=s=2$, we can consider $\lambda=1$ and $d=5$. The corresponding quantum code has parameters $[[273,255,4]]_{2}$, which improves the parameters $[[274,248,4]]_{2}$ and $[[273,246,4]]_{2}$ from [3]. If we consider $m=3$ instead, we obtain the parameters $[[4369,4337,4]]_{2}$, which surpasses the Gilbert-Varshamov bound from [9].